OpenStudy (anonymous):

x+6/x^2-3x=A/x+B/x-3 Find A-B

Parth (parthkohli):

Partial fractions madness!\[\rm {x + 6 \over x^2 - 3x} = {A \over x} + {B \over x - 3}\]

Parth (parthkohli):

The first thing to do is this:\[\rm {x+6 \over x(x - 3)} = {A \over x} + {B \over x - 3}\]

Parth (parthkohli):

Now, multiply by the common denominator which is \(\rm x(x - 3)\) (both sides).

Parth (parthkohli):

\[\rm x + 6 = A(x - 3) + B(x)\]

Parth (parthkohli):

Now, substitute \(\rm x = -3\) to get the value of \(\rm B\).

Parth (parthkohli):

And to get the value of \(\rm A\), substitute \(\rm x = 0\).

OpenStudy (anonymous):

so B = 3 and A = -2?

OpenStudy (anonymous):

A=-2 B=3

Parth (parthkohli):

No...

Parth (parthkohli):

Substitute \(\rm x = 3\) to get the value of \(\rm B\).

Parth (parthkohli):

Yes, so B = 3.

Parth (parthkohli):

Yeah, A = -2 B = 3

OpenStudy (anonymous):

okay thanks!!

Parth (parthkohli):

-2 - 3 = -5